# Improved Prediction Of Ligand-Protein Binding Affinities By Meta-Modeling Ho-Joon Lee1,2,*#, Prashant S. Emani3,*# and Mark B. Gerstein3,4,5,6,7#

### v2

@@ Acknowledgements
The authors report no competing interests.
@@
Selection of the receptor structure file: All data files for each complex in the PDBbind dataset (Liu et al., 2017) are contained in separate subfolders. We utilized the protein PDB file for the remaining processing steps.
@@
Preparation of the receptor structure: The receptor structure is read in as a PDB file and preprocessed in PyMol (https://pymol.org). We used the pymol class and associated functions within our pipeline script. The preprocessing consists of the removal of all solvent and heteroatoms in the molecules. The resulting PDB file is then converted into the requisite PDBQT format using OpenBabel (http://openbabel.org). We used the obabel flags ‘-e = continue after errors’, ‘-p = add hydrogens appropriate for the pH’, and ‘–partialcharge gasteiger’ to calculate Gasteiger partial charges. The pH was set at 7.0. We included the PDBQT-specific flags ‘r = output as a rigid molecular’ and ‘p = Preserve atom indices from input file’. The implementation of this is written as
@@
Creation of the configuration file: The configuration file serves to provide a set of commands to the smina package (Koes et al., 2013) (as well as the original AutoDock Vina tool (Trott and Olson, 2010)). The file contains the receptor PDBQT file, the ligand SDF file, the center coordinates of the search box in the receptor structure file coordinate system, and the dimensions of the search box. The calculation of the center coordinates varies depending on the datasets under consideration. For the PDBbind dataset, we used the center of mass of each ligand SDF file as the center of the search box, employing PyMol’s centerofmass function to do so. The search box was set to have dimensions of 40 Angstroms in the x, y and z directions.
@@
We used linear regression, ElasticNet, and LASSO algorithms for linear meta-models and the XGBoost algorithm for a non-linear meta-model implemented in scikit-learn in Python (Pedregosa et al., 2011). The steps in the preprocessing of the PDBbind dataset were as follows:
@@
To understand the contributions of ligand properties to performance of all the models, we used the web tool SwissADME (Daina et al., 2017) to extract the properties of the PDBbind CoreSet ligands from their SMILES strings. We first converted the ligand SMILES strings to canonical SMILES format using OpenBabel (http://openbabel.org). We then submitted the canonical SMILEs strings to SwissADME’s web interface. This yielded a set of 46 properties for 263 ligands in the CoreSet. We subsequently sought to measure the correlation between the deviation of the predicted scores from the experimental values with the numerical features, to identify those ligand features that might contribute to reduced performance of the model.
@@
To identify protein features that might be associated with better or worse performance of specific tools, we extracted the relevant subsets of PDBbind CoreSet PDB IDs from our analyses and fed them into the “Retrieve/ID mapping” tool of the UniProt database (www.uniprot.org/id-mapping) (The UniProt Consortium, 2023). The corresponding UniProt IDs and the following features were downloaded in a .tsv file: ’From’, ’Entry’, ’Reviewed’, ’Entry Name’, ’Protein names’, ’Gene Names’, ’Organism’, ’Length’, ’Active site’, ’Binding site’, ’Function [CC]’, ’Catalytic activity’, ’Annotation’, ’Gene Ontology (biological process)’, ’Helix’, ’Turn’, ’Beta strand’, ’Coiled coil’, ’Compositional bias’, ’Domain [CC]’, ’Domain [FT]’, ’Motif’, ’Protein families’, ’Region’, ’Repeat’, ’Zinc finger’, ’Pathway’, ’Gene Ontology (molecular function)’, ’Post-translational modification’. We first filtered out all UniProt entries that were not labeled “reviewed”. Next, we manually checked repeats, where the same PDB ID mapped to multiple UniProt IDs, by looking up the PDB IDs on the RCSB PDB website (www.rcsb.org/). We checked both the “Gene Names” and “Species” and removed all UniProt entries that did not match the PDB entry. In all cases, this left only a single PDB ID-UniProt ID mapping for each PDB ID. Additionally, to run the tests of over- or under-representation, we also deduplicated the UniProt IDs. That is, we removed all cases of proteins that were represented more than once, as the CoreSet has a large number of proteins that recur in the dataset. This led to a significant reduction in the number of proteins tested for UniProt feature prioritization from 266 to 60. Fisher’s exact test was run when comparing the number of occurrences of a feature in the subset versus the whole. The features tested are: the number of times a given UniProt term was annotated at all; the occurrence of Gene Ontology (Aleksander et al., 2023) terms, both the “Biological Process” and “Molecular Function” categories; and protein Domain and Pathway annotations.
@@
Aleksander, S.A., Balhoff, J., Carbon, S., Cherry, J.M., Drabkin, H.J., Ebert, D., Feuermann, M., Gaudet, P., Harris, N.L., Hill, D.P., et al. (2023). The Gene Ontology knowledgebase in 2023. Genetics 224.
Daina, A., Michielin, O., and Zoete, V. (2017). SwissADME: a free web tool to evaluate pharmacokinetics, drug-likeness and medicinal chemistry friendliness of small molecules. Scientific reports 7, 42717.
Koes, D.R., Baumgartner, M.P., and Camacho, C.J. (2013). Lessons learned in empirical scoring with smina from the CSAR 2011 benchmarking exercise. J Chem Inf Model 53, 1893-1904.
Liu, Z., Su, M., Han, L., Liu, J., Yang, Q., Li, Y., and Wang, R. (2017). Forging the Basis for Developing Protein–Ligand Interaction Scoring Functions. Accounts of Chemical Research 50, 302-309.
Pedregosa, F., Varoquaux, G., Gramfort, A., Michel, V., Thirion, B., Grisel, O., Blondel, M., Prettenhofer, P., Weiss, R., Dubourg, V., et al. (2011). Scikit-learn: Machine Learning in Python. J Mach Learn Res 12, 2825–2830.
The UniProt Consortium (2023). UniProt: the Universal Protein Knowledgebase in 2023. Nucleic acids research 51, D523-d531.
Trott, O., and Olson, A.J. (2010). AutoDock Vina: improving the speed and accuracy of docking with a new scoring function, efficient optimization, and multithreading. J Comput Chem 31, 455-461.
Supplementary Tables
Table S1. Dataset sizes for the empirical docking outputs after applying different RMSD filters
Table S2. Deep learning run times
Table S3. Absolute errors (deviations) of CoreSet predictions by the docking and deep learning models
Table S4. Meta-model performances version 1
Table S5. Meta-model performances version 2
Table S6. Full prediction results for CoreSet
Table S7. Base model score correlations with SwissADME features
Table S8. XGBoost meta-model score correlations with SwissADME features
Table S9. UniProt and SwissADME enrichment analyses for ligands and proteins
@@
Figure 1. Overview of our workflow
Figure 2. Performances of empirical scoring functions and consensus filtering
Figure 3. Performances of deep learning models
(A) Scatter plots of experimental binding affinities (x axis) and model predictions (y axis) for the full 5,305 complexes in the PDBbind RefinedSet (266 complexes in the CoreSet benchmark). Each plot shows predictions by the average of the best models or PC1 in terms of Pearson correlation coefficients for the CoreSet in each training strategy. (B) Box-whisker plots of Pearson correlation coefficients (left panel) and mean squared errors (MSE; right panel) for the CoreSet by a total of 22 PDBbind-finetuned (M1 and M2) and PDBbind-trained (M3) models and PC1 models of M1, M2, and M3 along with SMINA and Vinardo models. Predictions of the 22 models are mean values from 10 repeated 5-fold CVs. M1, M2, and M3 correspond to D1F, D2F, and D3 in (A).
Figure 4. Model comparison of CoreSet predictions
(A) Heatmap of Pearson correlations between all pairs of CoreSet predictions by the 42 models from the docking and deep learning tools. The dendrograms are results of hierarchical clustering with Euclidean distance and the complete linkage method. (B) Bar plot of Pearson correlations between the experimental binding affinity and predictions by the 42 models for the CoreSet complexes. The models are ordered by the correlation values. (C) Heatmap of absolute errors or deviations between the experimental binding affinity and predictions by the 42 models (columns) for each of the CoreSet complexes (rows). The dendrograms are as in (A). (D) Box-whisker plot of the data used in (B). The models are ordered by the median values.
Figure 5. Performances of meta-models
[…]
Figure 6. Prediction synergy and complementarity of docking and DL tools by meta-modeling
(A) Absolute errors (AEs) of the mean predictions for the CASF-2016 benchmark set (266 complexes) by the 4 meta-models with ED-A-P (open circles; ordered by AEs). The blue circles correspond to 104 complexes (39.1%) whose predictions by the DL tools (DAP|PC1 in Fig. 4D) are better than the mean predictions by the 4 meta-models or the 4 docking tools (i.e., lower AEs). The red circles correspond to 39 complexes (14.7%) whose mean predictions by the 4 docking tools are better than those by the meta-models or the DAP|PC1. (B-D) Scatter plots of AEs of the (mean) predictions for the CASF-2016 benchmark set by the DAP|PC1 vs. the 4 meta-models (B), by the 4 docking tools vs. the 4 meta-models (C), and by the 4 docking tools vs. the DAP|PC1 (D). The blue and red dots in (B) and (C) correspond to those in (A). In (D), the 48 complexes (18.0%) in red indicate better predictions (i.e., lower AEs) by the docking tools on average, whereas the remaining 218 complexes (82.0%) in blue indicate better predictions by the DAP|PC1.
Figure S6. Model comparison of non-core (training set) predictions.
Figure S7. Feature importance of meta-models.

### v3

@@ Acknowledgements
HL reports a consulting role at Guidepoint outside of this submitted work. The other authors report no competing interests.
@@
Improved Prediction Of Ligand-Protein Binding Affinities By Meta-Modeling
Ho-Joon Lee1,2,*#, Prashant S. Emani3,*# and Mark B. Gerstein3,4,5,6,7#
1Dept. of Genetics and 2Yale Center for Genome Analysis, Yale University, New Haven, CT 06510, USA 3Dept. of Molecular Biophysics & Biochemistry, Yale University, New Haven, CT 06520, USA 4Program in Computational Biology & Bioinformatics, 5Dept. of Computer Science, 6Dept. of Statistics & Data Science, and 7Dept. of Biomedical Informatics & Data Science, Yale University, New Haven, CT 06520, USA
* Equal contributions
# Correspondence: HL, ho-joon.lee@yale.edu; PSE, prashant.emani@yale.edu; MBG, pi@gersteinlab.org
@@
Selection of the receptor structure file: All data files for each complex in the PDBbind dataset1 are contained in separate subfolders. We utilized the protein PDB file for the remaining processing steps.
@@
Preparation of the receptor structure: The receptor structure is read in as a PDB file and preprocessed in PyMol (https://pymol.org). We used the pymol class and associated functions within our pipeline script. The preprocessing consists of the removal of all solvent and heteroatoms in the molecules. The resulting PDB file is then converted into the requisite PDBQT format using OpenBabel (http://openbabel.org). We used the obabel flags ‘-e = continue after errors’, ‘-p = add hydrogens appropriate for the pH’, and ‘–partialcharge gasteiger’ to calculate Gasteiger partial charges. The pH was set at 7.0. We included the PDBQT-specific flags ‘r = output as a rigid molecular’ and ‘p = Preserve atom indices from input file’. The implementation of this is written as:
@@
Creation of the configuration file: The configuration file serves to provide a set of commands to the smina package2 (as well as the original AutoDock Vina tool3). The file contains the receptor PDBQT file, the ligand SDF file, the center coordinates of the search box in the receptor structure file coordinate system, and the dimensions of the search box. The calculation of the center coordinates varies depending on the datasets under consideration. For the PDBbind dataset, we used the center of mass of each ligand SDF file as the center of the search box, employing PyMol’s centerofmass function to do so. The search box was set to have dimensions of 40 Angstroms in the x, y and z directions.
@@
We used linear regression, ElasticNet, and LASSO algorithms for linear meta-models and the XGBoost algorithm for a non-linear meta-model implemented in scikit-learn in Python4. The steps in the preprocessing of the PDBbind dataset were as follows:
@@
To understand the contributions of ligand properties to performance of all the models, we used the web tool SwissADME5 to extract the properties of the PDBbind CoreSet ligands from their SMILES strings. We first converted the ligand SMILES strings to canonical SMILES format using OpenBabel (http://openbabel.org). We then submitted the canonical SMILEs strings to SwissADME’s web interface. This yielded a set of 46 properties for 263 ligands in the CoreSet. We subsequently sought to measure the correlation between the deviation of the predicted scores from the experimental values with the numerical features, to identify those ligand features that might contribute to reduced performance of the model.
@@
To identify protein features that might be associated with better or worse performance of specific tools, we extracted the relevant subsets of PDBbind CoreSet PDB IDs from our analyses and fed them into the “Retrieve/ID mapping” tool of the UniProt database (www.uniprot.org/id-mapping)6. The corresponding UniProt IDs and the following features were downloaded in a .tsv file: ’From’, ’Entry’, ’Reviewed’, ’Entry Name’, ’Protein names’, ’Gene Names’, ’Organism’, ’Length’, ’Active site’, ’Binding site’, ’Function [CC]’, ’Catalytic activity’, ’Annotation’, ’Gene Ontology (biological process)’, ’Helix’, ’Turn’, ’Beta strand’, ’Coiled coil’, ’Compositional bias’, ’Domain [CC]’, ’Domain [FT]’, ’Motif’, ’Protein families’, ’Region’, ’Repeat’, ’Zinc finger’, ’Pathway’, ’Gene Ontology (molecular function)’, ’Post-translational modification’. We first filtered out all UniProt entries that were not labeled “reviewed”. Next, we manually checked repeats, where the same PDB ID mapped to multiple UniProt IDs, by looking up the PDB IDs on the RCSB PDB website (www.rcsb.org/). We checked both the “Gene Names” and “Species” and removed all UniProt entries that did not match the PDB entry. In all cases, this left only a single PDB ID-UniProt ID mapping for each PDB ID. Additionally, to run the tests of over- or under-representation, we also deduplicated the UniProt IDs. That is, we removed all cases of proteins that were represented more than once, as the CoreSet has a large number of proteins that recur in the dataset. This led to a significant reduction in the number of proteins tested for UniProt feature prioritization from 266 to 60. Fisher’s exact test was run when comparing the number of occurrences of a feature in the subset versus the whole. The features tested are: the number of times a given UniProt term was annotated at all; the occurrence of Gene Ontology7 terms, both the “Biological Process” and “Molecular Function” categories; and protein Domain and Pathway annotations.
LIT-PCBA Virtual Screening
For the LIT-PCBA virtual screening exercise, we utilized balanced sets of 10 active and 10 inactive ligands randomly drawn from the combined training and validation cohorts of the LIT_PCBA database8. To run the docking tools on the datasets, we generated three-dimensional SDF structures for each of the ligands from their SMILES strings using the following command:
obabel -:’<SMILES string>’ –gen2d -omol | obabel -imol –minimize –noh -omol | obabel -imol -h -e –gen3d -osdf -O ’<output SDF file>’
We also utilized the LIT-PCBA-provided PDB structures for the protein-ligand complex to configure the docking, by removing solvent molecules and other heteroatoms from the mol2 structure of the protein and then using the center-of-mass coordinates of the corresponding docked ligand in the PDB complex to define the docking search box. This facilitates a search for poses within the same binding pocket as the template complex provided by LIT-PCBA.
@@
1. Liu, Z.; Su, M.; Han, L.; Liu, J.; Yang, Q.; Li, Y.; Wang, R., Forging the Basis for Developing Protein–Ligand Interaction Scoring Functions. Accounts of Chemical Research 2017, 50 (2), 302-309. 2. Koes, D. R.; Baumgartner, M. P.; Camacho, C. J., Lessons learned in empirical scoring with smina from the CSAR 2011 benchmarking exercise. J Chem Inf Model 2013, 53 (8), 1893-904. 3. Trott, O.; Olson, A. J., AutoDock Vina: improving the speed and accuracy of docking with a new scoring function, efficient optimization, and multithreading. J Comput Chem 2010, 31 (2), 455-61. 4. Pedregosa, F.; Varoquaux, G.; Gramfort, A.; Michel, V.; Thirion, B.; Grisel, O.; Blondel, M.; Prettenhofer, P.; Weiss, R.; Dubourg, V.; Vanderplas, J.; Passos, A.; Cournapeau, D.; Brucher, M.; Perrot, M.; Duchesnay, É., Scikit-learn: Machine Learning in Python. J. Mach. Learn. Res. 2011, 12 (null), 2825–2830. 5. Daina, A.; Michielin, O.; Zoete, V., SwissADME: a free web tool to evaluate pharmacokinetics, drug-likeness and medicinal chemistry friendliness of small molecules. Scientific reports 2017, 7 (1), 42717. 6. The UniProt Consortium, UniProt: the Universal Protein Knowledgebase in 2023. Nucleic acids research 2023, 51 (D1), D523-d531. 7. Aleksander, S. A.; Balhoff, J.; Carbon, S.; Cherry, J. M.; Drabkin, H. J.; Ebert, D.; Feuermann, M.; Gaudet, P.; Harris, N. L.; Hill, D. P.; Lee, R.; Mi, H.; Moxon, S.; Mungall, C. J.; Muruganugan, A.; Mushayahama, T.; Sternberg, P. W.; Thomas, P. D.; Van Auken, K.; Ramsey, J.; Siegele, D. A.; Chisholm, R. L.; Fey, P.; Aspromonte, M. C.; Nugnes, M. V.; Quaglia, F.; Tosatto, S.; Giglio, M.; Nadendla, S.; Antonazzo, G.; Attrill, H.; Dos Santos, G.; Marygold, S.; Strelets, V.; Tabone, C. J.; Thurmond, J.; Zhou, P.; Ahmed, S. H.; Asanitthong, P.; Luna Buitrago, D.; Erdol, M. N.; Gage, M. C.; Ali Kadhum, M.; Li, K. Y. C.; Long, M.; Michalak, A.; Pesala, A.; Pritazahra, A.; Saverimuttu, S. C. C.; Su, R.; Thurlow, K. E.; Lovering, R. C.; Logie, C.; Oliferenko, S.; Blake, J.; Christie, K.; Corbani, L.; Dolan, M. E.; Drabkin, H. J.; Hill, D. P.; Ni, L.; Sitnikov, D.; Smith, C.; Cuzick, A.; Seager, J.; Cooper, L.; Elser, J.; Jaiswal, P.; Gupta, P.; Jaiswal, P.; Naithani, S.; Lera-Ramirez, M.; Rutherford, K.; Wood, V.; De Pons, J. L.; Dwinell, M. R.; Hayman, G. T.; Kaldunski, M. L.; Kwitek, A. E.; Laulederkind, S. J. F.; Tutaj, M. A.; Vedi, M.; Wang, S. J.; D’Eustachio, P.; Aimo, L.; Axelsen, K.; Bridge, A.; Hyka-Nouspikel, N.; Morgat, A.; Aleksander, S. A.; Cherry, J. M.; Engel, S. R.; Karra, K.; Miyasato, S. R.; Nash, R. S.; Skrzypek, M. S.; Weng, S.; Wong, E. D.; Bakker, E.; Berardini, T. Z.; Reiser, L.; Auchincloss, A.; Axelsen, K.; Argoud-Puy, G.; Blatter, M. C.; Boutet, E.; Breuza, L.; Bridge, A.; Casals-Casas, C.; Coudert, E.; Estreicher, A.; Livia Famiglietti, M.; Feuermann, M.; Gos, A.; Gruaz-Gumowski, N.; Hulo, C.; Hyka-Nouspikel, N.; Jungo, F.; Le Mercier, P.; Lieberherr, D.; Masson, P.; Morgat, A.; Pedruzzi, I.; Pourcel, L.; Poux, S.; Rivoire, C.; Sundaram, S.; Bateman, A.; Bowler-Barnett, E.; Bye, A. J. H.; Denny, P.; Ignatchenko, A.; Ishtiaq, R.; Lock, A.; Lussi, Y.; Magrane, M.; Martin, M. J.; Orchard, S.; Raposo, P.; Speretta, E.; Tyagi, N.; Warner, K.; Zaru, R.; Diehl, A. D.; Lee, R.; Chan, J.; Diamantakis, S.; Raciti, D.; Zarowiecki, M.; Fisher, M.; James-Zorn, C.; Ponferrada, V.; Zorn, A.; Ramachandran, S.; Ruzicka, L.; Westerfield, M., The Gene Ontology knowledgebase in 2023. Genetics 2023, 224 (1). 8. Tran-Nguyen, V.-K.; Jacquemard, C.; Rognan, D., LIT-PCBA: An Unbiased Data Set for Machine Learning and Virtual Screening. Journal of Chemical Information and Modeling 2020, 60 (9), 4263-4273.
@@
[margin-bottom=0.0063in,margin-top=0mm,margin-right=0mm,margin-left=0mm,border-style=none,padding=0mm]![[Uncaptioned image]](extracted/5604664/a0000-img017.png)